# What is a chemostat? Insights from hybrid dynamics and stochastic thermodynamics

### v1

@@ I Introduction
To do so, we use a large volume limit Ball et al. (2006); Menz et al. (2012); Crudu et al. (2012); Kang and Kurtz (2013); Anderson and Kurtz (2015); Winkelmann and Schütte (2017); Hufton, Lin, and Galla (2019) developed for hybrid CRNs (summarized in Sec. III) where high-abundant species, whose molecule numbers scale (linearly) with the volume, are mixed together with low-abundant species, whose molecule numbers do not scale with the volume. Crucially, this volume limit is based on the Wentzel–Kramers–Brillouin (WKB) ansatz, Risken (1989); Touchette (2009); Bressloff (2014); Touchette (2018); Qian and Ge (2021); Falasco and Esposito (2025) avoiding physical inconsistencies found in previous scalings. Horowitz (2015); Ceccato and Frezzato (2018) It leads to a hybrid dynamics. The low-abundant species are interconverted by reactions (called discrete reactions) still described in terms of stochastic events. The high-abundant species can be interconverted only by reactions (called continuous reactions) described in terms of deterministic events.
@@ III.2 Dynamics
The hybrid dynamics is derived (as summarized in Appendix A) by using a systematic perturbation analysis in the limit $\Omega\to\infty$ of the CME (3) with finite molecule numbers $\boldsymbol{n}_{x}$ and finite concentrations $[\boldsymbol{y}]\equiv(\dots,[y],\dots)$, called partial macroscopic limitMenz et al. (2012); Winkelmann and Schütte (2017) The procedure leads to a closed dynamical description only under the assumption that the (scaled) reaction rates
@@
$\partial_{t}[\boldsymbol{y}|\boldsymbol{n}_{x}]_{t}=\sum\limits_{\rho\in\mathcal{R}_{c}}\mathbf{S}_{y,\rho}\,r_{\rho}(\boldsymbol{n}_{x},[\boldsymbol{y}|\boldsymbol{n}_{x}]_{t})\,,$ (14)
where $[\boldsymbol{y}|\boldsymbol{n}_{x}]_{t}$ is the vector of the concentrations of the high-abundant species at time $t$, given the molecule numbers $\boldsymbol{n}_{x}$ of the low-abundant species. Physically, the RRE (14) determines a continuous evolution of the concentrations of the high-abundant species depending on the low-abundant species. For this reason, the $\mathcal{R}_{c}$ reactions are called continuous reactions. On the other hand, the resulting dynamics for the molecule numbers of the low-abundant species is, to leading order, a jump process described by the following CME
$\small\begin{split}\partial_{t}p_{t}(\boldsymbol{n}_{x})=\sum\limits_{\rho\in\mathcal{R}_{d}}\big{\{}&r_{\rho}(\boldsymbol{n}_{x}-\mathbf{S}_{x,\rho},[\boldsymbol{y}|\boldsymbol{n}_{x}-\mathbf{S}_{x,\rho}]_{t})p_{t}(\boldsymbol{n}_{x}-\mathbf{S}_{x,\rho})\\ &-r_{\rho}(\boldsymbol{n}_{x},[\boldsymbol{y}|\boldsymbol{n}_{x}]_{t})p_{t}(\boldsymbol{n}_{x})\}\,.\end{split}$ (15)
Physically, the CME (15) determines a discrete variations of the molecule numbers of the low-abundant species depending on the high-abundant species. For this reason, the $\mathcal{R}_{d}$ reactions are called discrete reactions. Note that $p_{t}(\boldsymbol{n}_{x})$ is the leading order approximation of the marginalized joint probability distribution $P_{t}(\boldsymbol{n}_{x},\boldsymbol{n}_{y})$, i.e., $p_{t}(\boldsymbol{n}_{x})\simeq\sum_{\boldsymbol{n}_{y}}P_{t}(\boldsymbol{n}_{x},\boldsymbol{n}_{y})$ for $\Omega\to\infty$ (see Appendix A).
In summary, the molecule numbers $\boldsymbol{n}_{x}$ evolve according to a jump process (described by the CME (15)) whose reaction rates depend on the concentrations of the high-abundant species. During the dwell time between one jump and the next one, the concentrations $[\boldsymbol{y}]$ of the high-abundant species evolve according to a deterministic process (described by RRE (14)) whose rates depend on the molecule number of the low-abundant species. Furthermore, since the reaction rates featuring the CME (15) and the RRE (14) are of the same order because of the scaling condition in Eq. (13), the molecule numbers $\boldsymbol{n}_{x}$ and the concentrations $[\boldsymbol{y}]$ evolve on the same time scale.
@@ IV.1 Mass-action kinetics
for all $\rho\in\mathcal{R}_{c}$. This crucially implies that only hybrid CRNs with discrete and continuous reactions satisfying the stoichiometric conditions in Eqs. (17) and (18) have a well-defined dynamical description (given in Sec. III.2) in the partial macroscopic limit.
@@ V.1 Local detailed balance condition
$\begin{split}\Delta_{\rho}g(n_{y})&=(\mu^{\circ}_{y}-\ln n_{s})S_{y,\rho}+\ln(n_{y}+S_{y,\rho})!-\ln n_{y}!\\ &\approx(\mu^{\circ}_{y}-\ln n_{s})S_{y,\rho}+(u\ln u-u)\Big{|}_{\Omega[y]}^{\Omega[y]+S_{y,\rho}}\\ &\approx(\mu^{\circ}_{y}-\ln n_{s})S_{y,\rho}+\ln(\Omega[y])S_{y,\rho}\\ &=\big{(}\mu^{\circ}_{y}+\ln([y]/[s])\big{)}S_{y,\rho}\equiv\mu([y])\,S_{y,\rho}\,,\end{split}$ (30)
by using Stirling’s formula, i.e., $\ln n_{y}!\approx n_{y}\ln n_{y}-n_{y}$, a Taylor expansion for $S_{y,\rho}/(\Omega[y])\ll 1$, and by identifying the chemical potential of the high-abundant species as $\mu([y])\equiv\mu^{\circ}_{y}+\ln([y]/[s])$.
@@
for continuous reactions $\rho\in\mathcal{R}_{c}$. In practice, the local detailed balance condition (27) of hybrid CRNs converges, in the partial macroscopic limit, to the local detailed balance condition (5) of open CRNs with the low- and high-abundant species playing the role of internal species and chemostats, respectively.
@@ VI.1 Michaelis-Menten
Figure 1: Dynamics and thermodynamics of the Michaelies-Menten CRN (38). (a) Mean molecule number of the enzyme $E$ (with $\braket{n_{E}}_{t=0}$, $\braket{n_{E}}_{t\to\infty}$, and $\braket{n_{E}}^{(0)}_{t\to\infty}$ denoting its initial value, its long-time value resulting from the CME (3) for the closed Michaelies-Menten CRN, and its long-time value resulting from the CME (15) for the hybrid Michaelies-Menten CRN, respectively). (b) Mean concentration of the substrate $S$ (with $\braket{n_{S}}_{t=0}$ and $\braket{n_{S}}_{t\to\infty}$ denoting the initial mean molecule number and its long-time value resulting from the CME (3) for the closed Michaelies-Menten CRN, respectively). (c) Mean entropy production rate $\braket{\dot{\Sigma}}_{t}$. The initial distribution of the closed Michaelies-Menten CRN is $P_{t=0}(n_{E},n_{C},n_{S},n_{P})=\delta(n_{E}-1)\delta(n_{C})\delta(n_{S}-[S]_{0}\Omega)\delta(n_{P}-[P]_{0}\Omega)$, the initial distribution of the hybrid Michaelies-Menten CRN is $p_{t=0}(n_{E},n_{C})=\delta(n_{E}-1)\delta(n_{C})$, with the initial concentrations (in arbitrary units) $[S]_{0}=2$ and $[P]_{0}=1$. The rate constants are (in arbitrary units) $k_{+1}=k_{+2}=1/k_{-1}=1/k_{-2}=e$.
@@ VI.2 Cyclic Michaelis-Menten
Figure 2: Dynamics and thermodynamics of the cyclic Michaelies-Menten CRN (41). (a) Mean molecule number of the enzyme $E$ (with $\braket{n_{E}}_{t=0}$ and $\braket{n_{E}}_{t\to\infty}$, denoting its initial value and its long-time value resulting from the CME (3) for the closed cyclic Michaelies-Menten CRN, respectively). (b) Mean concentration of the substrate $S$ (with $\braket{n_{S}}_{t=0}$ and $\braket{n_{S}}_{t\to\infty}$ denoting the initial mean molecule number and its long-time value resulting from the CME (3) for the closed cyclic Michaelies-Menten CRN, respectively). (c) Mean entropy production rate of the discrete reactions $\braket{\dot{\Sigma}_{d}}_{t}$. (d) Mean entropy production rate of the continuous reaction $\braket{\dot{\Sigma}_{c}}_{t}$. The initial distribution of the closed cyclic Michaelies-Menten CRN is $P_{t=0}(n_{E},n_{C},n_{S},n_{P})=\delta(n_{E}-1)\delta(n_{C})\delta(n_{S}-[S]_{0}\Omega)\delta(n_{P}-[P]_{0}\Omega)$, the initial distribution of the hybrid cyclic Michaelies-Menten CRN is $p_{t=0}(n_{E},n_{C})=\delta(n_{E}-1)\delta(n_{C})$, with the initial concentrations (in arbitrary units) $[S]_{0}=2$ and $[P]_{0}=1$. The rate constants are (in arbitrary units) $\log k_{1}=\log k_{2}=-\log k_{-1}=-\log k_{-2}=1/2$ and $\log k_{3}=-\log k_{-3}=1$.
@@ VII Discussion & Conclusion
Figure 3: Exemplary hybrid CRN whose stoichiometry satisfies the scaling conditions in Eq. (13). (a) Closed CRN with low-abundant species $(X_{1},X_{2},X_{3})$ and high-abundant species $(Y_{1},Y_{2},Y_{3},Y_{4})$. (b) Open CRN with internal species $(X_{1},X_{2},X_{3})$ and chemostats represented by Erlenmeyer flasks of different colors. In the partial macroscopic limit, the closed CRN (a) converges to the open one (b). Arrows represent the forward chemical reactions using hypergraph notation, while backward reactions are not illustrated for simplicity. Solid arrows represent discrete reactions (Eq. (11)) that are unimolecular in the low-abundant (resp. internal) species, while having arbitrary stoichiometry in the high-abundant species (resp. chemostats). Dashed arrows represent continuous reactions (Eq. (10)) that do not involve low-abundant (resp. internal) species and have arbitrary stoichiometry in the high-abundant species (resp. chemostats).
@@
While our work clearly demonstrates how chemostats can result from an underlying closed CRN, it does so for a subset of all possible stoichiometries. This is a direct implication of the large volume limit developed in Refs. Ball et al., 2006; Menz et al., 2012; Crudu et al., 2012; Kang and Kurtz, 2013; Anderson and Kurtz, 2015; Winkelmann and Schütte, 2017; Hufton, Lin, and Galla, 2019 which is based on a single-scale perturbation analysis. To investigate the emergence of chemostats in CRNs not satisfying the stoichiometric conditions in Eqs. (17) and (18), a multi-scale perturbation analysis needs to be developed and used. We leave this to further studies.

### v2

@@ I Introduction
To do so, we use a large volume limit Ball et al. (2006); Menz et al. (2012); Crudu et al. (2012); Kang and Kurtz (2013); Anderson and Kurtz (2015); Winkelmann and Schütte (2017); Hufton, Lin, and Galla (2019) developed for hybrid CRNs (summarized in Sec. III) where high-abundant species, whose molecule numbers scale (linearly) with the volume, are mixed together with low-abundant species, whose molecule numbers do not scale with the volume. Crucially, this volume limit is based on the Wentzel–Kramers–Brillouin (WKB) ansatz, Risken (1989); Touchette (2009); Bressloff (2014); Touchette (2018); Qian and Ge (2021); Falasco and Esposito (2025) avoiding physical inconsistencies found in previous scalings. Horowitz (2015); Ceccato and Frezzato (2018) It leads to a hybrid dynamics. The low-abundant species are interconverted by reactions (called discrete reactions) still described in terms of stochastic events. The high-abundant species can be interconverted only by reactions (called continuous reactions) described in terms of deterministic rate equations.
@@ III.2 Dynamics
The hybrid dynamics is derived (as summarized in Appendix A) by using a systematic (single time-scale) perturbation analysis in the limit $\Omega\to\infty$ of the CME (3) with finite molecule numbers $\boldsymbol{n}_{x}$ and finite concentrations $[\boldsymbol{y}]\equiv(\dots,[y],\dots)$, called partial macroscopic limitMenz et al. (2012); Winkelmann and Schütte (2017) The procedure leads to a closed dynamical description only under the assumption that the (scaled) reaction rates
@@
$\partial_{t}[\boldsymbol{y}|\boldsymbol{n}_{x}]_{t}=\sum\limits_{\rho\in\mathcal{R}_{c}}\mathbf{S}_{y,\rho}\,r_{\rho}(\boldsymbol{n}_{x},[\boldsymbol{y}|\boldsymbol{n}_{x}]_{t})\\ \,,$ (14)
where $[\boldsymbol{y}|\boldsymbol{n}_{x}]_{t}$ is the vector of the concentrations of the high-abundant species at time $t$, given the molecule numbers $\boldsymbol{n}_{x}$ of the low-abundant species, evolving from a given initial condition $[\boldsymbol{y}|\boldsymbol{n}_{x}]_{t=0}=[\boldsymbol{y}]_{0}$. Physically, the RRE (14) determines a continuous evolution of the concentrations of the high-abundant species depending on the low-abundant species. For this reason, the $\mathcal{R}_{c}$ reactions are called continuous reactions. On the other hand, the resulting dynamics for the molecule numbers of the low-abundant species is, to leading order, a jump process described by the following CME
$\small\begin{split}\partial_{t}p_{t}(\boldsymbol{n}_{x})=\sum\limits_{\rho\in\mathcal{R}_{d}}\big{\{}&r_{\rho}(\boldsymbol{n}_{x}-\mathbf{S}_{x,\rho},[\boldsymbol{y}|\boldsymbol{n}_{x}-\mathbf{S}_{x,\rho}]_{t})p_{t}(\boldsymbol{n}_{x}-\mathbf{S}_{x,\rho})\\ &-r_{\rho}(\boldsymbol{n}_{x},[\boldsymbol{y}|\boldsymbol{n}_{x}]_{t})p_{t}(\boldsymbol{n}_{x})\}\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}\end{split}$ (15)
where $p_{t}(\boldsymbol{n}_{x})$ is the probability of there being $\boldsymbol{n}_{x}$ molecules of the low-abundant species at time $t$, evolving from a given initial condition $p_{t=0}(\boldsymbol{n}_{x})$. Physically, the CME (15) determines a discrete variations of the molecule numbers of the low-abundant species depending on the high-abundant species. For this reason, the $\mathcal{R}_{d}$ reactions are called discrete reactions. Note that $p_{t}(\boldsymbol{n}_{x})$ is the leading order approximation of the marginalized joint probability distribution $P_{t}(\boldsymbol{n}_{x},\boldsymbol{n}_{y})$, i.e., $p_{t}(\boldsymbol{n}_{x})\simeq\sum_{\boldsymbol{n}_{y}}P_{t}(\boldsymbol{n}_{x},\boldsymbol{n}_{y})$ for $\Omega\to\infty$ (see Appendix A).
In summary, the molecule numbers $\boldsymbol{n}_{x}$ evolve according to a jump process (described by the CME (15)) whose reaction rates depend on the concentrations of the high-abundant species. During the dwell time between one jump and the next one, the concentrations $[\boldsymbol{y}]$ of the high-abundant species evolve according to a deterministic process (described by RRE (14)) whose rates depend on the molecule number of the low-abundant species. Furthermore, since the reaction rates featuring the CME (15) and the RRE (14) are of the same order because of the scaling condition in Eq. (13), the molecule numbers $\boldsymbol{n}_{x}$ and the concentrations $[\boldsymbol{y}]$ evolve on the same time scale. For this reason, the partial macroscopic limit corresponds to a single time-scale perturbation analysis that cannot resolve reactions occurring on multiple time-scales whose rates do not satisfy the scaling condition in Eq. (13).
@@ IV.1 Mass-action kinetics
for all $\rho\in\mathcal{R}_{c}$. This crucially implies that only hybrid CRNs with discrete and continuous reactions satisfying the stoichiometric conditions in Eqs. (17) and (18) (and illustrated in Fig. 3) have a well-defined dynamical description (given in Sec. III.2) in the partial macroscopic limit.
@@
Remark. When the stoichiometric conditions in Eqs. (17) and (18) are not satisfied, reaction occur on multiple time-scales (proportional to powers of $1/\Omega$) and, therefore, cannot be resolved with a single time-scale perturbation analysis such as the partial macroscopic limit.
@@ V.1 Local detailed balance condition
$\small\begin{split}\Delta_{\rho}g(n_{y})&=(\mu^{\circ}_{y}-\ln n_{s})S_{y,\rho}+\ln(n_{y}+S_{y,\rho})!-\ln n_{y}!\\ &\approx(\mu^{\circ}_{y}-\ln n_{s})S_{y,\rho}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Omega\{u\ln(\Omega u)-u\}\Big{|}_{[y]}^{[y]+S_{y,\rho}/\Omega}}\\ &\approx(\mu^{\circ}_{y}-\ln n_{s})S_{y,\rho}+\ln(\Omega[y])S_{y,\rho}\\ &=\big{(}\mu^{\circ}_{y}+\ln([y]/[s])\big{)}S_{y,\rho}\equiv\mu([y])\,S_{y,\rho}\,,\end{split}$ (30)
by using Stirling’s formula, i.e., $\ln n_{y}!\approx n_{y}\ln n_{y}-n_{y}$, for the first approximation, a Taylor expansion for ${S_{y,\rho}/\Omega\ll 1}$ for the second approximation, and by identifying the chemical potential of the high-abundant species as $\mu([y])\equiv\mu^{\circ}_{y}+\ln([y]/[s])$ in the final step.
@@
for continuous reactions $\rho\in\mathcal{R}_{c}$ due to the stoichiometric conditions in Eq. (18). In practice, the local detailed balance condition (27) of hybrid CRNs converges, in the partial macroscopic limit, to the local detailed balance condition (5) of open CRNs with the low- and high-abundant species playing the role of internal species and chemostats, respectively.
@@ VI.1 Michaelis-Menten
Figure 1: Dynamics and thermodynamics of the Michaelis-Menten CRN (38). (a) Mean molecule number of the enzyme $E$ (with $\braket{n_{E}}_{t=0}$, $\braket{n_{E}}_{t\to\infty}$, and $\braket{n_{E}}^{(0)}_{t\to\infty}$ denoting its initial value, its long-time value resulting from the CME (3) for the closed Michaelis-Menten CRN, and its long-time value resulting from the CME (15) for the hybrid Michaelis-Menten CRN, respectively). (b) Mean concentration of the substrate $S$ (with $\braket{n_{S}}_{t=0}$ and $\braket{n_{S}}_{t\to\infty}$ denoting the initial mean molecule number and its long-time value resulting from the CME (3) for the closed Michaelis-Menten CRN, respectively). (c) Mean entropy production rate $\braket{\dot{\Sigma}}_{t}$. The initial distribution of the closed Michaelis-Menten CRN is $P_{t=0}(n_{E},n_{C},n_{S},n_{P})=\delta(n_{E}-1)\delta(n_{C})\delta(n_{S}-[S]_{0}\Omega)\delta(n_{P}-[P]_{0}\Omega)$, the initial distribution of the hybrid Michaelis-Menten CRN is $p_{t=0}(n_{E},n_{C})=\delta(n_{E}-1)\delta(n_{C})$, with the initial concentrations (in arbitrary units) $[S]_{0}=2$ and $[P]_{0}=1$. The rate constants are (in arbitrary units) $\log k_{+1}=\log k_{+2}=-\log k_{-1}=-\log k_{-2}=1$.
@@ VI.2 Cyclic Michaelis-Menten
Figure 2: Dynamics and thermodynamics of the cyclic Michaelis-Menten CRN (41). (a) Mean molecule number of the enzyme $E$ (with $\braket{n_{E}}_{t=0}$ and $\braket{n_{E}}_{t\to\infty}$, denoting its initial value and its long-time value resulting from the CME (3) for the closed cyclic Michaelis-Menten CRN, respectively). (b) Mean concentration of the substrate $S$ (with $\braket{n_{S}}_{t=0}$ and $\braket{n_{S}}_{t\to\infty}$ denoting the initial mean molecule number and its long-time value resulting from the CME (3) for the closed cyclic Michaelis-Menten CRN, respectively). (c) Mean entropy production rate of the discrete reactions $\braket{\dot{\Sigma}_{d}}_{t}$. (d) Mean entropy production rate of the continuous reaction $\braket{\dot{\Sigma}_{c}}_{t}$. The initial distribution of the closed cyclic Michaelis-Menten CRN is $P_{t=0}(n_{E},n_{C},n_{S},n_{P})=\delta(n_{E}-1)\delta(n_{C})\delta(n_{S}-[S]_{0}\Omega)\delta(n_{P}-[P]_{0}\Omega)$, the initial distribution of the hybrid cyclic Michaelis-Menten CRN is $p_{t=0}(n_{E},n_{C})=\delta(n_{E}-1)\delta(n_{C})$, with the initial concentrations (in arbitrary units) $[S]_{0}=2$ and $[P]_{0}=1$. The rate constants are (in arbitrary units) $\log k_{1}=\log k_{2}=-\log k_{-1}=-\log k_{-2}=1/2$ and $\log k_{3}=-\log k_{-3}=1$.
@@ VII Discussion & Conclusion
Figure 3: Exemplary hybrid CRN whose stoichiometry satisfies the conditions in Eqs. (17) and (18). (a) Closed CRN with low-abundant species $(X_{1},X_{2},X_{3})$ and high-abundant species $(Y_{1},Y_{2},Y_{3},Y_{4})$. (b) Open CRN with internal species $(X_{1},X_{2},X_{3})$ and chemostats represented by Erlenmeyer flasks of different colors. In the partial macroscopic limit, the closed CRN (a) converges to the open one (b). Arrows represent the forward chemical reactions using hypergraph notation, while backward reactions are not illustrated for simplicity. Solid arrows represent discrete reactions (Eq. (11)) that are unimolecular in the low-abundant (resp. internal) species, while having arbitrary stoichiometry in the high-abundant species (resp. chemostats). Dashed arrows represent continuous reactions (Eq. (10)) that do not involve low-abundant (resp. internal) species and have arbitrary stoichiometry in the high-abundant species (resp. chemostats).
@@
While our work clearly demonstrates how chemostats can result from an underlying closed CRN, it does so for a subset of all possible stoichiometries. This limitation arises from the partial macroscopic limit developed in Refs. Ball et al., 2006; Menz et al., 2012; Crudu et al., 2012; Kang and Kurtz, 2013; Anderson and Kurtz, 2015; Winkelmann and Schütte, 2017; Hufton, Lin, and Galla, 2019 which is based on a single time-scale perturbation analysis (as discussed at the end of subs. III.2), while multi-molecular reactions evolve over multiple time-scales. Hence, to investigate the emergence of chemostats in CRNs not satisfying the stoichiometric conditions in Eqs. (17) and (18), a multiple time-scale perturbation analysis Kuehn (2015) needs to be first developed and then used. We leave this to further studies.